# Visual Alignment of Medical Vision-Language Models for Grounded Radiology Report Generation

### v1

@@ 5.3 Comparison with State-of-the-Art Methods
Figure 4: Generated responses with VALOR. Top Row: LLaVA-Med [li2023llava] fails to detect subtle reasoning patterns needed for accurate diagnosis, resulting in hallucination of disease observations such as ‘‘consolidation’’. In contrast, VALOR with verifiable guidance leads to significantly better reasoning and diagnosis capabilities. After Stage-2, VALOR’s visual enhanced reasoning leads to correctly diagonizing all observations showcasing the need for awareness of disease-relevant portions and sustained attention. Bottom Row: LLaVA-Med again hallucinates diseases due to insufficient visual focus, in comparison VALOR reduces hallucinations and generates reports faithful to the X-Ray exhibiting better anatomical and semantic awareness.
@@
Clinical Reasoning and Generation Quality. We highlight how VALOR can significantly generate more visually aligned and clinically enriched reports in Figure 4, w.r.t to the baseline Med-VLM (LLavA-Med [li2023llava]) and the ground truth report. Note that VALOR, despite not using any externally trained model or curated medical preferences, is still able to capture the disease-relevant portions of the X-ray (top row) and is able to correctly capture the disease observations (such as “atelectasis,” “cardiomegaly,” “edema,” etc.). In the second row, LLaVA-Med not only hallucinated disease observations that had no correlation with the input image but also generated generic text with no medical terminology. VALOR without the visual alignment performs significantly better than the base med-vlm but still fails in cases where the semantic knowledge of disease prevalent regions is needed. In contrast, VALOR generates reports with disease-relevant observations that closely align with the input X-ray, demonstrating superior clinical reasoning abilities.
Table 1: Detailed performance comparison on the report generation task covering IU-Xray and MIMIC-CXR datasets. BL denotes BLEU. We compare against the following baselines: LLaVa-Med [li2023llava], SFT, DPO [rafailov2023direct] w/ MMedPO [zhu2024mmedpo] and GRPO [shao2024deepseekmathpushinglimitsmathematical] on natural langauge generation quality and coherence.
Models IU-Xray MIMIC-CXR
BL-1 BL-2 BL-3 BL-4 ROUGE-L METEOR BL-1 BL-2 BL-3 BL-4 ROUGE-L METEOR
LLaVA-Medv1.5 22.29 7.95 2.90 1.05 7.58 4.56 9.62 3.66 1.48 0.63 9.42 5.62
+ SFT 38.42 13.40 4.74 1.67 10.31 10.95 29.41 10.19 3.58 1.26 9.38 7.71
+ DPO [rafailov2023direct] 41.63 15.13 5.56 2.03 12.95 17.13 29.61 10.29 3.61 1.27 9.45 7.81
+ MMedPO [zhu2024mmedpo] 55.58 23.93 10.36 4.40 29.52 34.16 31.37 10.93 4.26 1.52 10.96 9.88
+ GRPO [shao2024deepseekmathpushinglimitsmathematical] 59.02 26.50 11.80 5.10 31.80 35.90 34.10 11.80 4.25 1.45 10.30 9.65
@@
Table 2: CheXbert disease classification metrics (F1) on the IU-Xray dataset. Scores are computed using the CheXpert label set diseases; higher values indicate better alignment with expert disease labels.
Model Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$
LLaVA-Med v1.5 40.6 38.4 38.6
MMedPO [zhu2024mmedpo] 47.2 48.6 49.8
RADAR [hou-etal-2025-radar] 51.2 52.6 52.4
@@
Table 3: Zero-shot comparison for the report generation task on the IU-Xray [demner2015preparing] dataset against state-of-the-art multimodal medical and non-medical LLMs. Avg BL refers to the mean BLEU score.
Type Models Avg BL $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$
General VLM GPT-4o [hurst2024gpt] 15.32 18.80 11.50
@@
Medical VLM MAIRA-2 [bannur2024maira] 20.22 30.60 32.36
@@
Visual Reasoning and Grounding Abilities. We also visualize which portions of the image are being focused on by each baseline in Figure 5 by visualizing the gradient overlay of each model on the X-ray. Clearly, it can be seen that not only LLaVA-Med fails to focus on disease-relevant areas, but it also looks at regions that have no bearing on where the potential diseases can be, leading to inaccurate diagnoses, i.e. “no visible abnormalities.” Stage-1 fine-tuned VALOR performs relatively better and is able to detect the opacity in the right lung, but the greater focus on the left lung and pelvis region underlines the insufficient anatomical knowledge in the model. After performing visual alignment, the region where VALOR focuses on becomes much more concentrated and accurate, which is reflected in the generated report which has the correct observations in this context “opacity”, “atelectasis”, and “consolidation.”
Quantitative Analysis. We evalaute VALOR prmarily on two tasks: generation quality and clinical accuracy on the two radiology report benchmarks IU-XRay [demner2015preparing] and MIMIC-CXR [johnson2020mimic], outperforming state-of-the-art preference tuning method MMedPO [zhu2024mmedpo], DPO [rafailov2023direct] and VALOR with only Stage-1 alignment in Table 5.3. We obtain a significant improvement over MMedPO on all three natural language metrics and VALOR w/ visual alignment. To assess clinical accuracy, we further report CheXpert-based disease accuracy in Table 5.3. On IU-Xray, VALOR achieves the best F-1 score (57.2%), outperforming RADAR [hou-etal-2025-radar] (52.4%) and clearly improving over LLaVA-Med (38.6%) and MMedPO (49.8%). As it can be seen from the overall NLG metrics, the alignment step of our model is crucial for obtaining a large improvement on both the MIMIC-CXR and IU-Xray datasets. Overall, these results demonstrate that VALOR not only produces more fluent and semantically aligned reports but also maintains strong clinical accuracy, establishing a new state of the art despite not using any curated preference data, trained retrieval models, or databases.
@@ 5.4 Ablation Studies
Figure 5: Attention Map Visualization. For the input X-Ray $\mathcal{I}$, we visualize the specific areas the Med-VLM [li2023llava] and VALOR focuses on while generating the report. Bottom Left: is the attention map for LLaVA-Med [li2023llava] which misses the lung region due to lack of anatomical reasoning, leading to the generated report containing the inaccurate conclusion that no diseases can be detected. Upper Right: is the attention map with Stage-1 results of VALOR. It can be seen that it specifically learns to spot the clinically relevant regions for generating the report but still seems to focus on non-important regions at the base of the X-Ray. Bottom Right: In comparison, VALOR allocates more attention on the right lung region, the most important area for this particular X-Ray demonstrating better visual localization and reasoning capabilities. <report> tags are omitted here for clarity.
Comparison with multimodal LLMs. We also compare our model with mainstream proprietary multimodal LLMs such as GPT-4o [hurst2024gpt] and SOTA medical LLM MAIRA-2 [bannur2024maira] in a zero-shot setting in Table 5.3. Our method outperforms GPT-4o [hurst2024gpt] by 17% and MAIRA-2 [bannur2024maira] by an average margin of 5% across the NLG metrics. Therefore, VALOR consistently performs better than both general-purpose and medical multimodal LLMs, showing more anatomy of visual knowledge along with enhanced clinical accuracy despite the significantly larger pre-training knowledge corpus of these models on large amounts of data.
Effect of clinical and visual guidance. In Figure 6, on the left, we ablate the Stage-2 mixing weight $\lambda_{\mathrm{vis}}$ Eq. (10): BLEU-4, ROUGE-L, and CheXbert F1 peak for $\lambda_{\mathrm{vis}}\!\in\![0.4,0.6]$. Small values under-use visual evidence (mimicking Stage-1), whereas large values overweight similarity and erode fluency, indicating a balance between $R_{\text{text}}$ and the rank-normalized visual advantage. Figure 6 examines the clinical reward period $k_{\text{clin}}$: too frequent signals (small $k_{\text{clin}}$) destabilize training and bias toward terse mentions; too infrequent (large $k_{\text{clin}}$) under-leverages Stage 1 guidance. A moderate $k_{\text{clin}}$ yields the best trade-off, as shown in Figure 6. Overall, Stage-1 builds verifiable, clinically consistent text with periodic supervision, whereas Stage-2 strengthens visual grounding by mixing visual advantages at each policy optimization step.
Figure 6: Evolution of Clinical and visual periodic guidance. Left: VALOR performs better with an optimal value of $\lambda_{\mathrm{vis}}\sim 0.5$. Right: VALOR performs better with periodic injection of clinical rewards $R_{clin}$ per $k_{clin}$ steps instead of frequent (per-step) guidance which destabilizes performance.
@@ 6 Conclusion
We introduced VALOR, a novel multimodal alignment framework that enhances the clinical and visual reasoning capabilities of Medical VLMs. To this end, we propose an alignment framework which utilizes both textual and visual rewards for optimization thereby enhancing the model’s medical multimodal comprehension capabilities. Our results demonstrate the versatility of VALOR across a diverse set of datasets. Further our visualization studies show the effectiveness of our alignment approach, highlighting the crucial role the image plays in the overall task of radiology report generation. Our findings show VALOR’s potential to advance medical image understanding and clinical healthcare. Future work will include more ablations on reward shaping and curriculum schedules to clarify which components drive gains and to mitigate reward hacking. These directions chart a practical path toward image-faithful, clinically reliable report generation systems.

### v2

@@ 5.2 Comparison with State-of-the-Art Methods
Enhanced Clinical Reasoning and Generation Quality. We highlight how VALOR can significantly generate more visually aligned and clinically enriched reports in Fig. 4, w.r.t. to the baseline Med-VLM and the ground truth report. Note that VALOR, despite not using any externally trained model or curated expert preference data, is still able to capture the disease-relevant portions of the X-ray (top row) and is able to correctly capture the disease observations (such as “atelectasis”, “cardiomegaly”, “edema” etc.). In the second row, LLaVA-Med not only hallucinated disease observations that had no correlation with the input image but also generated generic text with no medical terminology showing suboptimal to no reasoning abilities at all. VALOR without the visual alignment performs significantly better than the base Med-VLM but still fails in cases where the semantic knowledge of disease prevalent regions is needed. In contrast, VALOR generates reports with disease-relevant observations that closely align with the input X-ray, demonstrating superior clinical reasoning abilities.
Figure 5: Attention Map Visualization. For the input X-Ray $\mathcal{I}$, we visualize the specific areas the Med-VLM [li2023llava] and VALOR focuses on while generating the report. Top Right, Both Cases: is the attention map for the MedVLM [li2023llava] which misses the lung region due to lack of anatomical reasoning, leading to the generated report containing the inaccurate conclusion that no diseases can be detected. Lower Right, Both Cases: is the attention map without visual guidance results of VALOR. It can be seen that it specifically learns to spot the clinically relevant regions for generating the report but still seems to focus on non-important regions at the base of the X-Ray. Bottom Left, Both Cases: In comparison, VALOR allocates more attention on the right lung region, the most important area for this X-Ray demonstrating better visual localization and reasoning capabilities. <report> tags are omitted here for clarity.
Visual Reasoning and Grounding Abilities. We also visualize which portions of the image are being focused on by each baseline in Fig. 5 by visualizing the attention tokens of each model on the X-ray. Clearly, it can be seen that not only the baseline fails to focus on disease-relevant areas, but it also looks at regions that have no bearing on where the potential diseases can be, leading to inaccurate diagnoses, i.e. “no visible abnormalities.” VALOR without the visual grounding proposed in Sec. 4.2 performs relatively better and is able to detect the opacity in the right lung, but the greater focus on the left lung and pelvis region underlines the insufficient anatomical knowledge in the model. After performing visual alignment, the region where VALOR focuses on becomes much more concentrated and accurate, which is reflected in the generated report which has the correct observations in this context “opacity”, “atelectasis”, and “consolidation”, encouraging the model to look before generation.
Table 1: We compare the performance of VALOR against the following baselines: base Med-VLM [li2023llava], SFT, naive DPO [rafailov2023direct], GRPO [shao2024deepseekmathpushinglimitsmathematical] and against SOTA post-trained model MMedPO [zhu2024mmedpo] on the generation quality [papineni2002bleu, lin2004rouge, denkowski2011meteor] of reports as well as clinical relevance and accuracy [smit2020chexbert, ostmeier2024green, zhao2024ratescore]. The effectiveness of visual alignment is seen in the last two rows in both benchmarks. Overall VALOR achieves the best performance on both benchmarks. The best results are in bold and the second best in italics.
@@
Comparison w.r.t. SOTA Multimodal LLMs. We also compare our model with mainstream proprietary multimodal LLMs such as GPT-4o [hurst2024gpt] and medical LLMs [moor2023med, bannur2024maira, chen2024huatuogpt] in a zero-shot setting in Table 5.2. Our method outperforms GPT-4 series [hurst2024gpt], Gemini [comanici2025gemini] by  +10% and SOTA medical MLLMs by an average margin of  +12% in generation quality. Therefore, VALOR consistently performs better than both general-purpose and medical multimodal LLMs, showing more anatomy of visual knowledge along with enhanced clinical accuracy despite the significantly larger pre-training knowledge corpus of these models on large amounts of data.
Table 2: We compare the performance of VALOR on both report generation benchmarks against proprietary SOTA medical and non-medical VLMs on generation quality of reports [papineni2002bleu, lin2004rouge, denkowski2011meteor]. VALOR significantly outperforms in both settings highlighting the effectiveness of generating visually aligned reports. The best results are highlighted in bold and the second best in italics.
Type Proprietary Models IU-Xray MIMIC-CXR
BLEU $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$ BLEU $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$
General VLM GPT-4o [hurst2024gpt] 15.32 18.80 11.50 12.10 17.20 10.30
GPT-4.1 [achiam2023gpt] 16.48 19.95 12.34 12.94 17.88 10.92
Gemini-2.5 Pro [comanici2025gemini] 17.26 21.10 13.05 13.42 18.76 11.48
@@
Medical VLM MedFlamingo [moor2023med] 14.85 24.10 22.44 11.76 21.95 19.84
MAIRA-2 [bannur2024maira] 20.22 30.60 32.36 18.40 28.90 30.10
HuatuoGPT-V-8B [chen2024huatuogpt] 18.76 28.40 29.05 17.12 27.35 28.44
@@ 5.3 Ablation Studies
Figure 6: Evolution of Clinical and visual periodic guidance. Left: VALOR performs best with an optimal value of $\lambda_{\mathrm{vis}}\sim 0.5$. Right: VALOR performs better with periodic injection of clinical rewards $R_{clin}$ per $k_{clin}$ steps instead of frequent (per-step) guidance which destabilizes performance.
Figure 7: Progression of generation quality with policy iterations. We plot how the performance of the generation quality progresses with each policy iteration during Stage-1 and Stage-2. As it can be particularly seen in the leftmost figure, there is a significant performance jump at the onset of the clinical guidance ($k_{clin}$) and at the beginning of Stage-2 highlighting the importance of our two-stage multimodal alignment mechanism.
Need for Multi-Stage Reward Guidance. Fig. 6(a) ablates the Stage-2 mixing weight $\lambda_{\mathrm{vis}}$ in Eq. (8). We observe that BLEU-4, ROUGE-L, and CheXbert F1 consistently peak when $\lambda_{\mathrm{vis}}\!\in\![0.4,0.6]$, indicating that neither purely textual nor purely visual optimization is sufficient. Lower weightage under-utilizes visual informativeness and reduces Stage-2 to Stage-1 behavior, whereas large values collapses training towards only disease terms. This confirms that gains of VALOR arise from a balanced combination of textual faithfulness and visual grounding, which is further reflected in improvements reported in Table 4.
Clinical/Visual Intervention. Fig. 6(b) studies the periodicity of clinical reward injection through $k_{\text{clin}}$. Very frequent supervision (small $k_{\text{clin}}$) destabilizes optimization and biases the policy toward short, overly conservative clinical phrases, while overly sparse supervision weakens the clinical signal and under-utilizes Stage-1 guidance. A moderate reward period yields the best trade-off, as also shown in Table 4. Fig. 7 justifies this design by showing clear performance jumps both at the onset of periodic clinical guidance and again at the beginning of Stage-2, demonstrating that clinical intervention improves report faithfulness early on, while visual intervention further strengthens grounded generation.
[…]
$\lambda_{\mathrm{vis}}$ IU-Xray MIMIC-CXR
$uparrow BLEU\uparrow F1\uparrow$
0.0 25.61 55.8 12.90 61.9
[…]
$k_{\mathrm{clin}}$ IU-Xray MIMIC-CXR
$uparrow BLEU\uparrow F1\uparrow$
[…]
rare 26.30 55.1 12.14 62.4
@@ 6 Conclusion
We introduced VALOR, a novel multimodal alignment framework that enhances the clinical and visual reasoning capabilities of existing Medical VLMs. To this end, we propose an alignment framework which utilizes both textual and visual rewards for optimization thereby enhancing the model’s clinical multimodal comprehension capabilities. Our results demonstrate the versatility of VALOR across a diverse set of datasets. Further our visualization studies show the effectiveness of our alignment approach, highlighting the crucial role the image plays in the overall task of radiology report generation. Our findings show VALOR’s potential to advance medical image understanding and clinical healthcare. Future work can include more holistic reward designs and curriculum learning to clarify which components drive gains. These directions chart a practical path toward image-aligned and clinically reliable report generation systems.